# Pressure, Reference Points, and Risk-Taking Behavior: Evidence from Bench Press Competitions

### v1

@@ 4.1 Choice of attempt weight
where $\tilde{W}_{it}^{k}$ is the difference between lifter $i$’s attempt weight compared to his current best in competition $t$ at a $k$-th attempt, $X_{it}$ is observed characteristics of lifter $i$ and competition $t$ including gender, body weight, number of competition experiences, a dummy variable for first participation, and fixed effects for competition $t$’s equipment category, age class, division, weight class, and federation, $Z_{it}^{k}$ is a pressure variable during attempt defined later, $\varepsilon_{it}$ is assumed to be an i.i.d. error, and $\beta$ and $\gamma$ are vectors of coefficient parameters of $X_{it}$ and $Z_{it}^{k}$. Large $\tilde{W}_{it}^{k}$ implies his choosing a more challenging attempt weight. Our primary interest lies in $\gamma$, that is, the sensitivity to pressure.
The pressure variables, $Z_{it}^{k}$, which are exogenous due to the sequential game setting, represent the difference between the lifter’s current realized outcome and the attempted weight of the lower-ranked competitor at the next attempt, and the difference between the lifter’s current realized outcome and the expected attempt weight of the higher-ranked competitor. These variables capture the potential for the lifter’s rank to change by overtaking or being overtaken by the rivals.
@@ 5.1 Choice of attempt weight
Pressure, lower rival, 2nd 0.264
Pressure, higher rival, 2nd 0.264
Pressure, lower rival, 3rd 0.502
Pressure, higher rival, 3rd -0.092
@@
Num.Obs. 267062 267062
R2 0.570 0.375
R2 Adj. 0.568 0.372
RMSE 34.46 41.87
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, and federation fixed effects. The standard errors in the brackets are clustered at federation categories. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from lower rival at 2nd is (lower rival 2nd attempt) - (his 1st outcome). Pressure from lower rival at 3rd is (lower rival 3rd attempt) - (his 2nd outcome). Pressure from higher rival at 2nd is (higher rival 2nd attempt) - (his 1st outcome). Pressure from higher rival at 3rd is (higher rival 3rd attempt) - (his 2nd outcome).
Table 6 presents the regression results for the difference between the second and third attempt weights compared to the current best, with respect to various pressure variables. Male lifters show mixed results, with a statistically significant negative coefficient for the second attempt (-1.140) and a positive coefficient for the third attempt (5.232). This suggests that male lifters are less likely to challenge in the second attempt but are more inclined to take risks in the third attempt. Body weight and experience are positively associated with higher attempt weights in both attempts, indicating that heavier and more experienced lifters tend to choose higher weights. The positive coefficients for first participation (33.022 and 34.819) indicate that first-time participants tend to choose significantly higher attempt weights compared to their current best, which is substituted by the achieved outcome in the first attempt.
For the second attempt, pressure from both lower and higher rivals shows a statistically significant positive effect (0.264 for both), suggesting that the presence of rivals encourages lifters to choose higher weights. In contrast, for the third attempt, the pressure from lower rivals has a positive association with increased weights (0.502), while pressure from higher rivals has a negative effect (-0.092). This negative coefficient implies that when the current score of a lifter is distant from that of a higher-ranked rival during the third attempt, the lifter tends to adopt a more conservative strategy by selecting a lower weight.
@@ IV X X
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, and federation. The standard errors in the brackets are clustered at federation categories.
Table 7 provides insights into how pressure influences the success probabilities of the second and third lifting attempts. In Column 1, the pressure from being turned around in the second attempt significantly increases the likelihood of success, as indicated by the strong positive coefficients (0.072 and 0.037). This suggests that lifters exert greater effort when facing the fear of loss aversion, possibly driven by the desire to avoid a lower rank. In contrast, Column 2 shows that the pressure of being turned around by lower-ranked lifters in the third attempt has a less pronounced effect on success, with a small positive but relatively low coefficient (0.009). This may imply that lifters exert less effort to maintain their rank in the third attempt, likely due to fatigue and the greater difficulty of improving marginal gains in ranking at this stage.
However, in both attempts, the pressure of turning around higher-ranked lifters increases the probability of success. This is evidenced by the significant positive coefficient for the third attempt (0.049), indicating that lifters are motivated to exert more effort to improve their rank when facing upward pressure. Columns 3 and 4, which present the results of the IV regression, show similar patterns. Overall, these findings highlight the shifting dynamics of pressure and motivation between the second and third attempts. Loss aversion appears to play a crucial role in the second attempt, while efforts to climb in rank are significant in both attempts, as anticipated in Table 5.
@@ 6.1 Attempt and success probability
Panel (a) in Figure 4 illustrates the distribution of the proportional change of second and third attempt weights relative to the current best from benchmark to counterfactual “no pressure during attempt” setting. In the second attempt, the attempt weight minus the current best weight under counterfactual is significantly lower than the benchmark. For example, a lifter whose current best weight is 100kg would attempt 100kg under no pressure but 101 kg under actual pressure, as shown in the peaks around -0.2 and -0.9.111111The example shows $((100-100)-(101-100))/100=-1$. On the other hand, the third attempt distribution is more spread out, with some lifters attempting weights significantly above their benchmark. This is consistent with balancing the opposite signs of estimated coefficients of pressure. This demonstrates that lifters are more cautious in their second attempts, while third attempts tend to be riskier, with more ambitious weights.
@@
Panel (b) in Figure 4 displays the distribution of the proportional change of second and third attempt success probabilities from benchmark to counterfactual. The changes in success probabilities are heavily concentrated below zero for both the second and third attempts, indicating that lifters would be more likely to fail under no pressure. The second attempt distribution shows a larger concentration in the range between -0.2 and 0.02. For example, a lifter who achieves the lift by 60% in the sense of objective probability would succeed in lifting by 51 % under no pressure, as shown in the peak around -0.15.121212The example shows $(51-60)/60=-0.15$. The third attempt distribution shifts more towards zero, which means that the effect of pressure is milder in the third attempt than the second attempt. Overall, success probabilities would decrease under no pressure from rivals.
@@ 6.2 Expected weight
Figure 5 presents the distribution of expected achieved weight differences from the lifters’ current best for second and third attempts, comparing the conditions of “no pressure” during the attempt, lifting, or both.
For the second attempt, the expected outcome reveals a concentration of attempts around -0.2 and -0.9, particularly in the “no pressure in both” setting, suggesting that lifters would tend to attempt weights significantly below their current best in the absence of pressure. In contrast, the “no pressure during lifting” setting displays a sharp peak near zero, indicating a large number of lifters attempting weights slightly below their personal best without pressure during the lift. These patterns suggest that, in the absence of pressure, the expected outcome for most lifters would be lower than the benchmark, largely due to less challenging attempts.
In the third attempt, the expected outcome differs markedly from the second attempt, with a greater concentration of attempts close to zero, particularly in the “no pressure in both” and “no pressure during lifting” settings. Lifters are more inclined to attempt weights near or slightly above their current best in the third attempt compared to the second. This suggests that, under no pressure, the expected outcome might exceed the benchmark for a significant portion of lifters, mainly due to more challenging attempts, albeit with a slight decrease in success probability.
As practical advice for individual lifters seeking to maximize their expected achieved weight, each lifter should consider adjusting their attempt plan based on the pressure experienced during the second attempt, while maintaining commitment to the original plan in the third attempt, disregarding pressure from rivals. This strategy is crucial for final rankings. Similarly, in broader contexts such as maximizing expected gains from risky choices, individuals should commit to their original risky decisions, ignoring external pressures from competitors.
@@ 7.1 Choice of lifter’s current best score
As our reference point, lifter’s current best, is crucial, we need to take care of the construction. For example, lifters might consider the last 12 months best score as the current best and ignore the more past best score because older personal bests may not accurately reflect lifters’ current abilities. Although we partly take care about the issues by controlling age-class, we need to check robustness of our reference point choice. We consider personal best over the more recent competition, that is, the personal best achieved within a certain recent period (e.g., the last 12 months).
@@
Pressure, lower rival, 2nd 0.299
(0.007) 1
Pressure, higher rival, 2nd 0.297
Pressure, lower rival, 3rd 0.551
Pressure, higher rival, 3rd -0.100
@@
Num.Obs. 248778 248778
R2 0.628 0.367
R2 Adj. 0.627 0.364
RMSE 31.09 40.89
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, and federation fixed effects. The standard errors in the brackets are clustered at federation categories. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from lower rival at 2nd is (lower rival 2nd attempt) - (his 1st outcome). Pressure from lower rival at 3rd is (lower rival 3rd attempt) - (his 2nd outcome). Pressure from higher rival at 2nd is (higher rival 2nd attempt) - (his 1st outcome). Pressure from higher rival at 3rd is (higher rival 3rd attempt) - (his 2nd outcome).
@@ IV X X
Num.Obs. 248778 248778 248778 248778
R2 0.047 0.020 0.044 0.020
R2 Adj. 0.042 0.016 0.040 0.016
RMSE 0.45 0.49 0.45 0.49
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, and federation. The standard errors in the brackets are clustered at federation categories.
Table 8 corresponds to Table 6 for the analysis of attempt weights using the lifter’s current best within one year. The main results remain robust under this alternative specification. Similarly, Table 9 corresponds to Table 7 for the analysis of success probabilities, utilizing the lifter’s current best within a year. Most of the main results are consistent when using this alternative measure. The only notable difference is that the coefficient for the dummy variable indicating “turning around the higher-ranked rival” shows a statistically significant positive effect.

### v2

@@ 4.1 Choice of attempt weight
where $\tilde{W}_{it}^{k}$ is the difference between lifter $i$’s attempt weight compared to his current best in competition $t$ at a $k$-th attempt, $X_{it}$ is observed characteristics of lifter $i$ and competition $t$ including gender, body weight, number of competition experiences, a dummy variable for first participation, and fixed effects for competition $t$’s equipment category, age class, division, weight class, and federation. It also includes a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor, potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules, and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. $Z_{it}^{k}$ is a pressure variable during attempt defined later, $\varepsilon_{it}$ is assumed to be an i.i.d. error, and $\beta$ and $\gamma$ are vectors of coefficient parameters of $X_{it}$ and $Z_{it}^{k}$. Large $\tilde{W}_{it}^{k}$ implies his choosing a more challenging attempt weight. Our primary interest lies in $\gamma$, that is, the sensitivity to pressure.
The pressure variables, $Z_{it}^{k}$, which are exogenous due to the sequential game setting, represent the difference between the lifter’s current realized outcome and the attempted weight of the lower-ranked competitor at the next attempt, and the difference between the lifter’s current realized outcome and the expected attempt weight of the higher-ranked competitor predicted by the linear regression model. The actual higher-ranked attempt weight should not be used because it is not realized at lifter $i$’s attempt, although it must be predictable. We discuss the predictability issues and how to correct prediction errors, in particular, for constructing standard errors via bootstrap in Section 7. These variables capture the potential for the lifter’s rank to change by overtaking or being overtaken by the rivals.
@@ 5.1 Choice of attempt weight
Pressure, lower rival, 2nd 0.106
Pressure, higher rival, 2nd 0.449
Pressure, lower rival, 3rd 0.057
Pressure, higher rival, 3rd 0.409
@@
Num.Obs. 246521 246521
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. The standard errors in the brackets are clustered at federation categories. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from lower rival at 2nd is (lower rival 2nd attempt) - (his 1st outcome). Pressure from lower rival at 3rd is (lower rival 3rd attempt) - (his 2nd outcome). Pressure from higher rival at 2nd is (higher rival 2nd attempt) - (his 1st outcome). Pressure from higher rival at 3rd is (higher rival 3rd attempt) - (his 2nd outcome).
Table 5 presents the regression results for the difference between the second and third attempt weights compared to the current best, incorporating various pressure variables. Male lifters exhibit mixed responses, with a statistically significant negative coefficient for the second attempt (–1.279) and a positive coefficient for the third attempt (3.583). This suggests that male lifters adopt a conservative approach in their second attempt but take greater risks in their third attempt, a pattern consistent with the literature on gender differences in risk-taking. Body weight and experience are positively associated with higher attempt weights in both attempts, indicating that heavier and more experienced lifters generally select heavier weights. The positive coefficients for first participation (32.122 and 33.623) suggest that first-time competitors attempt significantly higher weights relative to their current best.
For the second attempt, pressure from both lower- and higher-ranked rivals has a statistically significant positive effect (0.106 and 0.449, respectively), suggesting that the presence of competitive pressure encourages lifters to attempt heavier weights. In the third attempt, pressure from lower-ranked rivals continues to have a positive effect (0.057), though its magnitude is smaller than in the second attempt. Meanwhile, pressure from higher-ranked rivals remains strongly positive (0.409), reinforcing the idea that lifters attempt heavier weights when their higher-ranked competitors perform well. The persistence of positive coefficients in the third attempt suggests that lifters under strong competitive pressure, particularly from higher-ranked rivals, do not necessarily adopt a conservative strategy but instead continue to take risks, likely in an effort to improve their final ranking.
@@ IV X X
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. The standard errors in the brackets are clustered at federation categories.
Table 6 provides insights into how pressure influences the success probabilities of the second and third lifting attempts. In Column 1, the pressure from being turned around in the second attempt has a small but positive effect on success probability (0.014), suggesting that lifters may exert additional effort to avoid losing rank. However, the coefficient is modest, indicating a limited influence of loss aversion in this context. Conversely, the pressure of turning around a higher-ranked lifter in the second attempt has a slightly negative coefficient (–0.012), implying that lifters attempting to overtake a higher rival may face increased difficulty in execution.
For the third attempt, the pressure of being turned around remains positively associated with success probability (0.033), while the pressure of turning around a higher-ranked rival has a negative coefficient (–0.038). This suggests that lifters are more likely to succeed when pressured from below but face greater challenges when attempting to surpass a higher-ranked competitor. Columns 3 and 4, which present the results of the IV regression, show consistent patterns with slightly reduced magnitudes and statistical significance.
Overall, these findings highlight the evolving role of pressure across attempts, where loss aversion plays a role, and the difficulty of overtaking higher-ranked rivals remains a significant challenge.
5.3 Heterogeneity
Tables 7 and 8 show how responses to pressure vary by gender, experience, and historical rivalry frequency. Male lifters respond more strongly to both lower- and higher-ranked rivals (0.104 and 0.455, respectively) than female lifters (0.068 and 0.257), particularly under upward pressure. More experienced lifters are more responsive to higher-ranked rivals (0.031–0.032), reinforcing the role of competitive cues accumulated through experience. A higher historical rivalry frequency, defined as the number of prior encounters with a rival, is associated with greater responsiveness (0.100 and 0.095 for second attempts; 0.127 and 0.095 for third attempts), suggesting that familiarity with rivals plays a role in shaping attempt selection. These trends indicate that male experienced lifters facing well-known rivals take greater risks.
For success probabilities, gender differences vary depending on the direction of pressure. When lifters are under pressure from lower-ranked rivals—i.e., at risk of being overtaken—male lifters consistently show higher success probabilities in both the second (0.015) and third attempts (0.032). Female lifters, by contrast, show little to no benefit in the second attempt (–0.003), but exhibit a positive response in the third attempt (0.034), suggesting that the effect of downward pressure converges across genders in later stages.
In contrast, when attempting to overtake higher-ranked rivals, both male and female lifters experience reduced success probabilities, with effects slightly more negative for female lifters (–0.039 vs. –0.034 in the third attempt). These results indicate that upward pressure imposes psychological and performance costs for most lifters, with gender gaps persisting, while the motivational effects of downward pressure appear more stage-dependent and gender-convergent.
Experience plays a limited role in moderating these effects, with small but statistically significant coefficients. Historical rivalry frequency is associated with higher success probabilities when under pressure from lower-ranked rivals (0.012 in the second attempt and 0.028 in the third), but with lower success probability when facing pressure from higher-ranked rivals (–0.007 in the third attempt). This pattern suggests that familiarity with lower-ranked rivals may help lifters manage performance anxiety, while familiarity with higher-ranked rivals may instead heighten psychological stress. These results emphasize the complex interplay between gender, experience, and competitive history in shaping both risk-taking behavior and execution under pressure.
Table 7: Attempt Heterogeneity on Pressure
(1) (2) (3) (4) (5) (6)
Dependent Variable $W_{it}^{2}-$(best) $W_{it}^{3}-$(best)
(Pressure, lower rival)$\times$ 1(female) 0.068 -0.021
(0.015) (0.026)
(Pressure, lower rival)$\times$ 1(male) 0.104 0.056
(0.009) (0.006)
(Pressure, higher rival)$\times$ 1(female) 0.257 0.154
(0.010) (0.015)
(Pressure, higher rival)$\times$ 1(male) 0.455 0.417
(0.010) (0.011)
(Pressure, lower rival)$\times$ (Num experience) 0.004 0.001
(0.001) (0.001)
(Pressure, higher rival)$\times$ (Num experience) 0.031 0.032
(0.001) (0.002)
(Pressure, lower rival)$\times$ (Historical rivalry freq) 0.100 0.127
(0.011) (0.010)
(Pressure, higher rival)$\times$ (Historical rivalry freq) 0.095 0.095
(0.012) (0.015)
Control X X X X X X
Num.Obs. 247449 247449 247449 247438 247438 247438
R2 0.535 0.400 0.332 0.346 0.292 0.256
R2 Adj. 0.533 0.397 0.329 0.343 0.289 0.253
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking as in the main specification. The standard errors in the brackets are clustered at federation categories.
Table 8: Success Probability Heterogeneity on Pressure
(1) (2) (3) (4) (5) (6)
Dependent Variable 1(success 2nd) 1(success 3rd)
1(Turned around)$\times$ 1(female) -0.003 0.034
(0.006) (0.013)
1(Turned around)$\times$ 1(male) 0.015 0.032
(0.003) (0.004)
1(Turning around)$\times$ 1(female) -0.014 -0.039
(0.007) (0.010)
1(Turning around)$\times$ 1(male) -0.003 -0.034
(0.006) (0.006)
1(Turned around)$\times$ (Num experience) 0.001 0.003
(0.000) (0.001)
1(Turning around)$\times$ (Num experience) 0.001 -0.002
(0.000) (0.000)
1(Turned around)$\times$ (Historical rivalry freq) 0.012 0.028
(0.003) (0.007)
1(Turning around)$\times$ (Historical rivalry freq) -0.002 -0.007
(0.004) (0.004)
Control X X X X X X
IV X X X X X X
Num.Obs. 246521 246521 246521 246521 246521 246521
R2 0.046 0.046 0.046 0.022 0.021 0.021
R2 Adj. 0.042 0.041 0.041 0.017 0.016 0.016
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking as in the main specification. The standard errors in the brackets are clustered at federation categories.
@@ 6.1 Attempt and success probability
Panel (a) in Figure 4 shows the distribution of proportional changes in second and third attempt weights relative to the current best under the counterfactual without pressure during the attempt. In the second attempt, most lifters select lighter weights than in the benchmark, resulting in a left-skewed distribution. In the third attempt, the distribution is more symmetric and concentrated near zero, with modest variation in both directions. These patterns suggest that pressure has a stronger and more uniform effect on second attempts than third attempts.
@@
Panel (b) in Figure 4 presents the distribution of proportional changes in second and third attempt success probabilities from the benchmark to the counterfactual. For the second attempt, the distribution is relatively concentrated, with peaks around zero, indicating that some lifters experience modest gains or losses in success probability when pressure is removed. In contrast, the third attempt shows a more dispersed distribution with heavier tails, suggesting greater variability in how lifters respond to the absence of pressure. While a subset of lifters benefit, many experience reduced success. Overall, the removal of competitive pressure leads to mixed effects in both attempts, with more pronounced heterogeneity in the third.
@@ 6.2 Expected weight
Notes: We use estimated coefficients in Table 5 and Columns (3) and (4) in Table 6. Positive values indicate better performance under no pressure than under actual competition pressure.
Figure 5 presents the distribution of expected achieved weight differences from the lifters’ current best for second and third attempts, comparing the conditions of “no pressure” during the attempt, lifting, or both. We take the differences to incorporate individual-level heterogeneity on the best achieved weight.
In the second attempt, all counterfactual settings lead to lower expected outcomes, especially in the “no pressure in both” condition, where the distribution is clearly left-skewed. The “no pressure during lifting” case shows a sharper peak near zero, suggesting that many lifters would still choose similar weights without pressure during execution.
In the third attempt, the distributions become more dispersed, particularly under “no pressure during attempt” and “no pressure in both” conditions. A small group under “no pressure during lifting” shows gains, but most lifters experience reduced outcomes. These patterns highlight substantial heterogeneity in pressure responses and reinforce the motivational role of real-time competition.
As practical advice, most lifters may benefit from adjusting their second attempts in response to competitive pressure, as removing pressure tends to reduce both attempted weight and success probability. In contrast, maintaining one’s original plan in the third attempt—rather than reacting to pressure from rivals—may help preserve focus, though this strategy is not universally optimal across lifters. More broadly, individuals in sequential high-stakes settings should account for their own sensitivity to competitive pressure, recognizing that for many, performance deteriorates in its absence.
@@ 7.1 Choice of lifter’s current best score
As our main outcome includes the lifter’s current best, we need to take care of the construction. For example, lifters might consider the last 12 months best score as the current best and ignore the more past best score because older personal bests may not accurately reflect lifters’ current abilities. Although we partly take care about the issues by controlling age-class, we need to check robustness of our reference point choice. We consider personal best over the more recent competition, that is, the personal best achieved within a certain recent period (e.g., the last 12 months).
@@
Pressure, lower rival, 2nd 0.092
Pressure, higher rival, 2nd 0.483
Pressure, lower rival, 3rd 0.043
Pressure, higher rival, 3rd 0.427
@@
Num.Obs. 230744 230744
R2 0.561 0.325
R2 Adj. 0.559 0.322
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. The standard errors in the brackets are clustered at federation categories. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from lower rival at 2nd is (lower rival 2nd attempt) - (his 1st outcome). Pressure from lower rival at 3rd is (lower rival 3rd attempt) - (his 2nd outcome). Pressure from higher rival at 2nd is (higher rival 2nd attempt) - (his 1st outcome). Pressure from higher rival at 3rd is (higher rival 3rd attempt) - (his 2nd outcome).
@@ IV X X
Num.Obs. 230744 230744 230744 230744
R2 0.043 0.022 0.042 0.021
R2 Adj. 0.038 0.017 0.037 0.016
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. The standard errors in the brackets are clustered at federation categories.
Table 9 corresponds to Table 5 for the analysis of attempt weights using the lifter’s current best within one year. The main results remain robust under this alternative specification. Male lifters continue to exhibit stronger responses in attempt weight selection, with coefficients of 0.528 statistically insignificant for the second attempt and 5.785 for the third attempt. Body weight remains positively associated with attempt weights (0.151 and 0.254), while experience shows a small negative effect (–0.059 and –0.125), suggesting that more experienced lifters may exhibit slightly more conservative choices.
The coefficients for pressure from lower- and higher-ranked rivals remain positive and statistically significant. For the second attempt, the effects are 0.092 (lower rival) and 0.483 (higher rival), while for the third attempt, the corresponding coefficients are 0.043 and 0.427. These patterns reinforce the idea that lifters are sensitive to competitive dynamics, particularly when higher-ranked rivals lift successfully, leading to upward adjustments in attempt weights.
Similarly, Table 10 corresponds to Table 6 for the analysis of success probabilities, utilizing the lifter’s current best within a year. The main results are largely consistent with the primary specification. Upward pressure—especially attempts to overtake higher-ranked rivals—is associated with a higher likelihood of failure.
Overall, the alternative specification using the current best within a year confirms the key findings from the main analysis: competitive pressure affects both strategic weight selection and execution success, and these effects persist across time horizons.